# Domain Adaptation based Object Detection for Autonomous Driving in Foggy and Rainy Weather

### v3

@@ Abstract
Typically, object detection methods for autonomous driving that rely on supervised learning make the assumption of a consistent feature distribution between the training and testing data, this such assumption may fail in different weather conditions. Due to the domain gap, a detection model trained under clear weather may not perform well in foggy and rainy conditions. Overcoming detection bottlenecks in foggy and rainy weather is a real challenge for autonomous vehicles deployed in the wild. To bridge the domain gap and improve the performance of object detection in foggy and rainy weather, this paper presents a novel framework for domain-adaptive object detection. The adaptations at both the image-level and object-level are intended to minimize the differences in image style and object appearance between domains. Furthermore, in order to improve the model’s performance on challenging examples, we introduce a novel adversarial gradient reversal layer that conducts adversarial mining on difficult instances in addition to domain adaptation. Additionally, we suggest generating an auxiliary domain through data augmentation to enforce a new domain-level metric regularization. Experimental findings on public V2V benchmark exhibit a substantial enhancement in object detection specifically for foggy and rainy driving scenarios.
@@
The objective of this paper is to reduce the domain gap between various weathers for enhanced object detection. To handle the domain shift problem (e.g. ClearFoggy and ClearRainy), in this paper, we present a new domain adaptation framework that aims to enhance the robustness of object detection in foggy and rainy weather conditions. Our proposed framework follows an unsupervised setting, similar to previous works [23, 24, 25]. In this setting, we have well-labeled clear-weather images as the source domain, while the foggy and rainy weather images, which serve as the target domains, lack any annotations. This unsupervised setting is because adverse weather images with labeling (manual annotating) are time-consuming and costly. Inspired by [23, 26], the proposed method aims to reduce the domain feature discrepancies in both image style and object appearance. To enhance robustness and prevent data-level overfitting, we propose a Dynamic Masking Process to generate masked images by “dropping” some pixel regions. To achieve domain-invariant features, we incorporate both image-level and object-level domain classifiers as components to facilitate domain adaptation in our CNN architecture. These classifiers are responsible for distinguishing between different domains. By employing an adversarial approach, our detection model learns to generate features that are invariant to domain variations, thereby confusing the domain classifiers. This adversarial design encourages the network to produce features that are agnostic to the specific weather conditions, leading to improved object detection performance in foggy and rainy weather scenarios.
@@ II-C Object Detection via Domain Adaptation
Domain adaptation is effective in reducing the distribution discrepancy between different domains, enabling models trained on a labeled source domain to be applicable to an unlabeled target domain. There has been a growing interest in addressing domain adaptation for object detection [23, 49, 50, 51, 52, 53, 54, 26] in recent years. Several studies [23, 55, 50, 54, 26] have explored the alignment of features from different domains to achieve DA object detectors. A DA Faster R-CNN framework [23] was proposed to reduces the domain gap at both the image level and instance level. He et al. [55] proposed a multi-adversarial network that aligns domain features and proposal features hierarchically to minimize domain distribution disparity. In addition to feature alignment, image style transfer approaches [34, 49, 56, 57] are utilized to address the challenge of DA. An image translation module [34] was utilized to convert images from the source domain to the target domain. They then trained the object detector using adversarial training on the target domain. [56] adopted a progressive image translation strategy and introduced a weighted task loss during adversarial training to address image quality differences. Several previous methods [58, 59, 60, 61] have also proposed complex architectures for domain adaptation in object detection. Feature Pyramid Networks (FPN) was utilized to incorporate pixel-level and category-level adaptation for object detection [58]. In order to incorporate the uncertainty of unlabeled target data, [60] introduced an uncertainty-guided self-training mechanism, which leverages a Probabilistic Teacher and Focal Loss. Different with these methods, our approach does not introduce additional learnable parameters to the Faster R-CNN. Instead, we utilize an AdvGRL and a Domain-level Metric Regularization based on triplet loss. A key difference between our method and previous domain adaptation approaches lies in the treatment of training samples. While existing methods often assume that training samples are at the same challenging level, our approach introduces the AdvGRL for adversarial hard example mining, specifically targeting the improvement of transfer learning performance. Additionally, to mitigate overfitting and improve domain adaptation, an auxiliary domain is generated and incorporated to domain-level metric regularization.
@@ III-D Object-level based Adaptation
Besides the global differences at the image level, objects within different domains may exhibit variations in terms of appearance, size, color, and other characteristics. To address this, Each region proposal generated by the ROI Pooling layer is considered as a potential object of interest. After obtaining the object-level domain representation via ROI pooling, we introduce an object-level domain classifier to discern the origin of the local features, which is implemented by three fully-connected layers. The objective of the object-level domain classifier is to align the distribution of object-level features across different domains. Similar to the image-level domain classifier, we utilize the BCE loss to train our object-level domain classifier:
@@ IV-B Experimental Setting
Dataset setting: We conducted two main experiments in this paper: 1) Clear to Foggy Adaptation, denoted as Clear CityscapesFoggy Cityscapes, the labeled training set of Clear Cityscapes [21] and the unlabeled training set of Foggy Cityscapes [22] are used as the source and target domains during training, respectively. Subsequently, the trained model was evaluated Foggy Cityscapes validation set to report the performance. Rainy Cityscapes training set is used as the Auxiliary Domain in this Clear to Foggy Adaptation experiment. 2) Clear to Rainy Adaptation, denoted as the Clear CityscapesRainy Cityscapes, where the labeled training set of Clear Cityscapes [21] and the unlabeled training set of Rainy Cityscapes are used as the source and target domains during training, respectively. Then the trained model was evaluated on Rainy Cityscapes validation set to report the performance. Foggy Cityscapes training set is used as the Auxiliary Domain in this Clear to Rainy Adaptation experiment. Additionally, we analyzed the transfer learning performance on different intensity levels of fog and rain (small, medium, and large).
Training setting: We utilize ResNet-50 as the backbone for the Faster R-CNN [18]. Following in [23, 18], during training, We utilize back-propagation and stochastic gradient descent (SGD) to optimize all the deep learning methods in our approach. The initial learning rate of for iterations is used in all model training. Afterward, the learning rate is reduced to and training continues for an additional iterations. Weight decay is set as and momentum is set as for all experiments. Each training batch consists of three images from the source, target, and auxiliary domains respectively. For comparison purposes, we set the value in the original GRL (Equation (4)) to . In the AdvGRL (Equation (5)), the hardness threshold is set to , which is computed by averaging the parameters in Equation (1) with setting ( and ). In the subsequent analysis, we refer to “Reg + AdvGRL” as our proposed DA method. Additionally, “Reg + AdvGRL + DMP” is designated as our enhanced DA method, termed DA+, which is named as “Ours+” for short in the tables.
Evaluation metrics: We calculate the Average Precision for each category and the mean Average Precision across all categories using an Intersection over Union threshold of .

### v4

@@ Abstract
Typically, object detection methods for autonomous driving that rely on supervised learning make the assumption of a consistent feature distribution between the training and testing data, this such assumption may fail in different weather conditions. Due to the domain gap, a detection model trained under clear weather may not perform well in foggy and rainy conditions. Overcoming detection bottlenecks in foggy and rainy weather is a real challenge for autonomous vehicles deployed in the wild. To bridge the domain gap and improve the performance of object detection in foggy and rainy weather, this paper presents a novel framework for domain-adaptive object detection. The adaptations at both the image-level and object-level are intended to minimize the differences in image style and object appearance between domains. Furthermore, in order to improve the model’s performance on challenging examples, we introduce a novel adversarial gradient reversal layer that conducts adversarial mining on difficult instances in addition to domain adaptation. Additionally, we suggest generating an auxiliary domain through data augmentation to enforce a new domain-level metric regularization. Experimental findings on public benchmark exhibit a substantial enhancement in object detection specifically for foggy and rainy driving scenarios. The code is available at https://github.com/jinlong17/DA-Detect.
@@
The objective of this paper is to reduce the domain gap between various weathers for enhanced object detection. To handle the domain shift problem (e.g. ClearFoggy and ClearRainy), in this paper, we present a new domain adaptation framework that aims to enhance the robustness of object detection in foggy and rainy weather conditions. Our proposed framework follows an unsupervised setting, similar to previous works [23, 24, 25]. In this setting, we have well-labeled clear-weather images as the source domain, while the foggy and rainy weather images, which serve as the target domains, lack any annotations. This unsupervised setting is because adverse weather images with labeling (manual annotating) are time-consuming and costly. Inspired by [23, 26], the proposed method aims to reduce the domain feature discrepancies in both image style and object appearance. To enhance robustness and prevent data-level overfitting, we propose a Dynamic Masking Process to generate masked images by “dropping” some pixel regions. To achieve domain-invariant features, we incorporate both image-level and object-level domain classifiers as components to facilitate domain adaptation in our CNN architecture. These classifiers are responsible for distinguishing between different domains. By employing an adversarial approach, our detection model learns to generate features that are invariant to domain variations, thereby confusing the domain classifiers. This adversarial design encourages the network to produce features that are agnostic to specific weather conditions, leading to improved object detection performance in foggy and rainy weather scenarios.
@@ II-C Object Detection via Domain Adaptation
Domain adaptation is effective in reducing the distribution discrepancy between different domains, enabling models trained on a labeled source domain to be applicable to an unlabeled target domain. There has been a growing interest in addressing domain adaptation for object detection [23, 49, 50, 51, 52, 53, 54, 26] in recent years. Several studies [23, 55, 50, 54, 26] have explored the alignment of features from different domains to achieve DA object detectors. A DA Faster R-CNN framework [23] was proposed to reduce the domain gap at both the image level and instance level. He et al. [55] proposed a multi-adversarial network that aligns domain features and proposal features hierarchically to minimize domain distribution disparity. In addition to feature alignment, image style transfer approaches [34, 49, 56, 57] are utilized to address the challenge of DA. An image translation module [34] was utilized to convert images from the source domain to the target domain. They then trained the object detector using adversarial training on the target domain. [56] adopted a progressive image translation strategy and introduced a weighted task loss during adversarial training to address image quality differences. Several previous methods [58, 59, 60, 61] have also proposed complex architectures for domain adaptation in object detection. Feature Pyramid Networks (FPN) was utilized to incorporate pixel-level and category-level adaptation for object detection [58]. In order to incorporate the uncertainty of unlabeled target data, [60] introduced an uncertainty-guided self-training mechanism, which leverages a Probabilistic Teacher and Focal Loss. Different with these methods, our approach does not introduce additional learnable parameters to the Faster R-CNN. Instead, we utilize an AdvGRL and a Domain-level Metric Regularization based on triplet loss. A key difference between our method and previous domain adaptation approaches lies in the treatment of training samples. While existing methods often assume that training samples are at the same challenging level, our approach introduces the AdvGRL for adversarial hard example mining, specifically targeting the improvement of transfer learning performance. Additionally, to mitigate overfitting and improve domain adaptation, an auxiliary domain is generated and incorporated into domain-level metric regularization.
@@ III-D Object-level based Adaptation
Besides the global differences at the image level, objects within different domains may exhibit variations in terms of appearance, size, color, and other characteristics. To address this, Each region proposal generated by the ROI Pooling layer is considered as a potential object of interest. After obtaining the object-level domain representation via ROI pooling, we introduce an object-level domain classifier to discern the origin of the local features, which is implemented by three fully connected layers. The objective of the object-level domain classifier is to align the distribution of object-level features across different domains. Similar to the image-level domain classifier, we utilize the BCE loss to train our object-level domain classifier:
@@ IV-B Experimental Setting
Dataset setting: We conducted two main experiments in this paper: 1) Clear to Foggy Adaptation, denoted as Clear CityscapesFoggy Cityscapes, the labeled training set of Clear Cityscapes [21] and the unlabeled training set of Foggy Cityscapes [22] are used as the source and target domains during training, respectively. Subsequently, the trained model was evaluated by the Foggy Cityscapes validation set to report the performance. Rainy Cityscapes training set is used as the Auxiliary Domain in this Clear to Foggy Adaptation experiment. 2) Clear to Rainy Adaptation, denoted as the Clear CityscapesRainy Cityscapes, where the labeled training set of Clear Cityscapes [21] and the unlabeled training set of Rainy Cityscapes are used as the source and target domains during training, respectively. Then the trained model was evaluated on Rainy Cityscapes validation set to report the performance. Foggy Cityscapes training set is used as the Auxiliary Domain in this Clear to Rainy Adaptation experiment. Additionally, we analyzed the transfer learning performance on different intensity levels of fog and rain (small, medium, and large).
Training setting: We utilize ResNet-50 as the backbone for the Faster R-CNN [18]. Following in [23, 18], during training, We utilize back-propagation and stochastic gradient descent (SGD) to optimize all the deep learning methods in our approach. The initial learning rate of for iterations is used in all model training. Afterward, the learning rate is reduced to and training continues for an additional iterations. Weight decay is set as and momentum is set as for all experiments. Each training batch consists of three images from the source, target, and auxiliary domains respectively. For comparison purposes, we set the value in the original GRL (Equation (4)) to . In the AdvGRL (Equation (5)), the hardness threshold is set to , which is computed by averaging the parameters in Equation (1) with setting ( and ). In the subsequent analysis, we refer to “Reg + AdvGRL” as our proposed DA method. Additionally, “Reg + AdvGRL + DMP” is designated as our enhanced DA method, termed DA+, which is named “Ours+” for short in the tables.
Evaluation metrics: We calculate the Average Precision for each category and the mean Average Precision across all categories using an Intersection over the Union threshold of .